OpenStudy (anonymous):

MEDALS WILL BE GIVEN HELP x(x-7)=60 and also (1/2)x^2 minus (7/6)x =1 SOLVE BY FACTORING PLEAS MEDALS LL BE GIVEN

OpenStudy (anonymous):

Solve both by factoring plz!!

OpenStudy (cwrw238):

x(x-7)=60 x^2 - 7x - 60 = 0 yes - you can solve this by factoring you need 2 number whose product is -60 and whose sum is - 7

OpenStudy (anonymous):

can you do # 2?

OpenStudy (cwrw238):

(1/2)x^2 - (7/6)x - 1 = 0 multiply thru by 6 3x^2 - 7x - 6 = 0

OpenStudy (cwrw238):

yes you can also solve this by factoring

OpenStudy (raden):

alternative (No. 1): x(x-7)=60 x (x-7) =12 * 5 identically both sides, the value of x already be solved :)

OpenStudy (anonymous):

thanks a ton!!! i was confused!

OpenStudy (cwrw238):

yw